OpenStudy (jpes2193):

(8-6i)(-4-4i) Help Please

OpenStudy (bloomlocke367):

do you have to multiply that?

OpenStudy (jpes2193):

If (a+bi)(a-bi)= a^2 +b^ Yes we have to multiply

OpenStudy (bloomlocke367):

you can use the foil method.

OpenStudy (jpes2193):

Right when I used the foil method on a problem similar to this I got it wrong.

OpenStudy (bloomlocke367):

are you sure you did it correctly?

OpenStudy (jpes2193):

well that one was 4i(-2-8i) so Im not 100% sure

OpenStudy (bloomlocke367):



OpenStudy (jpes2193):

32-32i

OpenStudy (bloomlocke367):

and you should have gotten 32-8i for the other one.. but anyways, close, but not quite... it's -32-32i

OpenStudy (jpes2193):

okay. and then when you do the -6*-4 half you get -24-24i

OpenStudy (bloomlocke367):

no.... because it's -6i... do you know i represents?

OpenStudy (jpes2193):

\[\sqrt{-1}\]

OpenStudy (bloomlocke367):

yep, so, you'd get|dw:1426523043013:dw

OpenStudy (bloomlocke367):

woah.. okay, hang on

OpenStudy (bloomlocke367):